Suppose a stock had an initial price of $87 per share, paid a dividend of $2.15 per share during the year, and had an ending share price of $98. 


a. Compute the percentage total return. (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

b. What was the dividend yield? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 

c. What was the capital gains yield? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) 
Answer #1

% total return=(End value-Beginning value+Dividend)/Beginning value

=(98-87+2.15)/87

=15.11%(Approx).

Dividend yield=Dividend/Beginning value

=2.15/87

=2.47%(Approx).

Capital gains yield=(End value-Beginning value)/Beginning value

=(98-87)/87

=12.64%(Approx).